I'm struggling on this one & it's probably easy but I just can't seem to get it, PLEASE HELP...
Quadratic Eq.
Jane invested some amount of money at a rate of 12% simple interest and some other amount at the rate of 10% simple interest. She received yearly interest of $130.Randy also invested in the same scheme, but he interchanged amounts invested and received $4 more interest. How much did each of them invest at different rates?

Answer by checkley71(8403)

.12X+.10Y=130 MULTIPLY BY -.10 (-.012X-.010Y=-13)
.10X+.12Y=134 MULTIPLY BY .12 (.012X+.0144Y=16.08)
NOW ADD:
.012X+.0144Y=16.08
-.012X-.010Y=-13
-----------------------------------------
.0044Y=3.08
Y=3.08/.0044
Y=700 INVESTED BY JANE @ 10%
.12X+.10*700=130
.12X+70=130
.12X=130-70
.12X=60
X=60/.12
X=500 INVESTED BY JANE @ 12%
PROOF USING RANDY'S INVESTMENTS
.12*700+.10*500=134
84+50=134
134=134